OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -x + 2y = 8 -x + 10y = 48

OpenStudy (anonymous):

subtract them to get -8y = -40 y = 5 x = 2

OpenStudy (anonymous):

(2) -(1) gives : 8y = 40 hence y = 5 sub y = 5 back to (1) get x = 2

OpenStudy (anonymous):

I need steps I am completely confused right now

OpenStudy (anonymous):

Minus the first equation from the second: -x + 2y = 8 - -x + 10y = 48 __________________ -8y = -40 -8y = -40 8y = 40 y = 5 Then put 5 back into an equation as your y value: -x + 2(5) = 8 -x + 10 = 8 -x = -2 x = 2

OpenStudy (anonymous):

second equation from first I mean

OpenStudy (anonymous):

thank you that's a little easier